A body is released from rest at the top of a plane inclined at 30° to the horizontal and 4.0m high.
If the coefficient of the friction between the body and the plane is 0.3 , calculate the time the body takes to reach the bottom of the plane
1
Expert's answer
2020-01-31T10:37:13-0500
"l=\frac{h}{\sin{30}}=\frac{4}{\sin{30}}=8\ m"

"F=ma=mg\sin{30}-\mu mg\cos{30}"

"a=g(\sin{30}-\mu \cos{30})=\frac{2l}{t^2}"

"9.8(\sin{30}-0.3 \cos{30})=\frac{2(8)}{t^2}"

"t=2.6\ s"